1. Consider the power sets Consider the w -1) 2 5n 1 (+1). Determine the interval of convence 2. Given that = for IE(-1,1
Answer #1

00 n =0 (1²4) gennal feom ga power series Cm (-a) for the giun dentes cuising (947) (n = (ighsn (121) >comparting with geneR- lim (1) x 50 x ((1+1) + 1) / insa (nu) (-1) s(14) - m o vie son camps 41) | (n*+)9%2C1)e mht!) - im In C+080) - n°(*) (20 bgium take no (xe (-1,1)) OL 4+2x 4( 1+ 2x) 4 (172) given lala 1 <1 .. The guen denies can be written as. tx which is oquabut we already showed n = 0 n =0 + cy) nakane) n=0 20 2x4 n=0 an or - 1 ) A. n. 36-1) & n=0 an on 2 (-1)+n.gen-1) n = 0 -1